OpenStudy (anonymous):

What is the equation of the line meeting the given conditions. Please put the equation in standard form. Containing E(4, 3) and F(6, 1)

OpenStudy (anonymous):

use the equation y=kx+b. put the points in the equation

OpenStudy (anonymous):

ok but it needs to be in y-y1=m(x-x1) form

OpenStudy (anonymous):

do you know the formula for the slope between two points?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

hmm formula is \[m=\frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (anonymous):

Hi, there is a two point formula in straight lines. \[ y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} \cdot (x - x_1) \] Substitute x1= 4, y1 = 3 x2 = 6, y2 = 1 and get the equation. \[ y - 3 = \frac{1 - 3}{6 - 4} (x - 4) \] \[ y - 3 = \frac{-2}{2} (x - 4) \] \[ y - 3 = -1 (x - 4) \] \[ y + - 3 + x - 4 = 0 \] \[ x + y - 7 = 0 \]

OpenStudy (anonymous):

where (x1,y1)=(4, 3) and (x2,y2)=(6, 1)

OpenStudy (anonymous):

wow someone just gave you the answer. lucky